A stone released at the top of the shaft of a well hits the water 1.55 s later. (a) How deep is the well? (b) At what speed does the stone hit the water?
Answer #1

answer)a) from kinematics we have

h=1/2gt2

h=1/2*9.8*1.552

h=11.8 m

so the depth is 11.8 m or 12 m

b) again from kinematics we have

v=u+at

here initial velocity u=0 ( since we start from rest)

v=at

v=9.8*1.55 s=15.2 m/s

so the speed is 15.2 m/s or 15 m/s